Two spherical cantaloupes of the same kind are sold at a fruit and vegetable stand. The circumference of one is 60 cm and that of the other os 50 cm. The larger melon is 1.5 times as expensive as the smaller. Which mellon is the better buy?
Found 2 solutions by Alan3354, robertb:
Answer by Alan3354(69443)

60 = 1.2 * 50
The circumference is a linear function of the radius.
The volume increases as the cube of the radius.
--> 1.2 times the circumference = 1.2^3 times the volume
1.2^3 = 1.728
--------------
The volume of the larger melon is 1.728 times the smaller.
It costs 1.5 times as much, so it's cheaper per unit volume.
How much more is seeds and rind is not stated, tho, and "better" is not a well defined parameter.

Answer by robertb(5830)

It's a toss up, I think. The bigger melon is 1.5 times more expensive than the smaller one, but you only get around 1.152 times more melon content (volume-wise) per buck. So what do you think?